# Newcastle-upon-Tyne, UK

### v4

@@ 1 Introduction
  • Ultimately, we advise against relying on commodity sensors as primary sources of randomness for security-critical applications, both on a single- and multiple-sensor basis.
The rest of this paper is organised as such: §2 discusses sensor-based security mechanisms and established entropy metrics. §3 explains our experiment design for entropy estimation, including the threat model and dataset selection. §4 presents empirical results of our analyses and §5 discusses the implications for system design. We conclude in §6 with recommendations for further work. Our analysis work is released publicly to foster future research. [^1] https://github.com/cgshep/entropy-collapse-mobile-sensors
@@ 2.1 Mobile Sensor-based Security Applications
The notion of ‘hardness’ is typically inferred using model evaluation metrics; it is not grounded in standard methods used in the field of authentication, such as Shannon [21] and min-entropy [22]. A general approach follows one whereby sensor data is collected from $N$ users from which various features are extracted in the time or frequency domain (e.g. cross-correlation, spectral energy, Hamming, Euclidean and mean-absolute distances) [4, 13, 3, 2]. Features are subsequently classified using simple threshold-based or supervised classification models, e.g. Support Vector Machine (SVM) and Random Forest. Calculating false positive (FPR) and negative (FNR) rates [1], precision and recall [8], EERs [14, 13], Receiver-Operator Curves (ROC) [11], and accuracy [11] are then used to evaluate the model with respect to distinguishing between legitimate and illegitimate samples. (User studies have also been employed to evaluate the effectiveness of sensor-based authentication systems [18, 1].)
We observe that the vast majority of proposals do not directly measure entropy, which is a crucial factor in determining whether sensor data cannot be easily predicted by an adversary. Some work has attempted to assess the entropy of sensor signals, but this represents a minority. T2Pair by Li et al. [23] yielded 32.3–38.5 bits of Shannon entropy, with a refinement by Wu et al. [24] reporting 51–54 bits of Shannon entropy. Even in the best cases, this is low relative to modern cryptographic standards. [^2] See AIS 20/31 [25]: FCS_RNG.1 specifies $\geq$240 and $\geq$250 bits of min- and Shannon entropy respectively for the effective internal state of a random number generator. Lastly, we point to adjacent work that has questioned the utility of mobile sensors. Markantonakis et al. [4], building on Gurulian et al. [5] and Shepherd et al. [6], presented a reproducibility study disputing the claimed effectiveness of mobile sensors when deployed under real-world time constraints for NFC-based contactless transactions. Using a 500ms window for data acquisition, as stipulated by the EMV payment protocol, 0.179–0.246 EER was reported depending on the given sensor combination. Sensors were thus deemed unsuitable for proximity and relay attack detection without posing usability and security issues in practice.
@@ 2.2 Sensor Entropy Analyses
In earlier work, Voris et al. [22] investigated the use of accelerometers as true random number generators (TRNGs), specifically aboard a WISP RFID tag and a Nokia N97 mobile phone. The authors find that min-entropy—defined in §2.3—is proportional to the motion applied to the device, and that stationary movement has the lowest min-entropy. Intrinsic noise from the sensor’s circuitry and seismic noise, and the sampling rate of its analog-to-digital converter (ADC) are considered other significant influences on the generated entropy. Min-entropy values of 3.1–11.4 bits were measured depending on the movement of the accelerometer. Lv et al. [26] analyse three mobile sensors on an undisclosed Xiaomi Redmi smartphone: a triaxial accelerometer, gyroscope, and magnetometer. Min-entropy values of 0.593–5.876 are reported, depending on the modality and the entropy estimation method used. Krhovják et al. [7] examine the entropy of image and audio data collected from mobile phone cameras and microphones respectively. Using Nokia N73 and E-Ten X500 and M700 phones, Shannon entropy values of 2.9 (microphone) and 2.408–5.376 (camera) are reported, with min-entropy of 0.5 (microphone) and 0.754–3.928 (camera). Hennebert et al. [27] present an analysis of 10 sensor modalities available on two wireless sensor monitors: a TI eZ430-RF2500 and a Zolertia Z1. The work uses a single-sensor analysis, presenting min-entropy values between 0–7.85. Motion sensors, e.g. accelerometer and vibration sensors, yielded the highest entropy.
@@
Mobile sensors have also been proposed as entropy sources in low-cost RNG designs. Suciu et al. [28] propose using a phone’s GPS module along with its accelerometer, gyroscope and orientation sensors. Using data from an HTC Google Nexus One, the approach passes the tests established in the NIST SP 800-22 [29] suite, but no precise entropy values are presented. Wallace et al. [30] explore the use of accelerometer, gyroscope, microphone, WiFi, GPS and camera data as randomness sources. Mobile sensors are inherently opportunistic entropy sources, contrasting with dedicated hardware like ring oscillators, those build on Johnson-Nyquist thermal noise, or quantum-based generators (e.g. see [31]). Mobile sensors depend heavily on user behaviour and environmental, introducing biases and correlations absent in controlled entropy sources. Existing sensor-based security mechanisms largely overlook these dynamics, relying instead on heuristic or model evaluation metrics [8, 11]. Such approaches do not account well for skewed distributions and other biases in the underlying data. This contrasts with typical entropy measures used in RNG analyses or in the study of authentication systems [32]. For instance, NIST SP800-90B [33] and AIS 20/31 [25] recommend the use of min-entropy to estimate ‘worst-case’ unpredictability of a given source. Our study bridges the gap by systematically evaluating entropy across modalities and datasets.
@@ 5.1 Comparison with Related Work
Prior studies have claimed that mobile sensors can serve as a robust source of security or randomness. Much work has relied on machine-learning-based metrics as a proxy to evaluate such claims [13, 14, 4, 2], with a smaller subset using more established entropy metrics. Li et al.[23] report 32–38 bits of Shannon entropy for some motion-based secure pairing schemes, and Wu et al.[24] estimate 51–54 bits. However, dedicated entropy analyses frequently demonstrate that individual sensors confer very little min-entropy, especially under worst-case assumptions [26, 27, 22, 7]. Our results, derived from a systematic analysis of multiple datasets and modalities (see Tables 25), corroborate and challenge these claims.

### v5

@@ 1 Introduction
  • Ultimately, we advise against relying on commodity sensors as sources of unpredictability for security-critical applications, both on a single- and multiple-sensor basis.
The rest of this paper is organised in the following way: §2 discusses sensor-based security mechanisms and established entropy metrics. §3 explains our experiment design for entropy estimation, including the threat model and dataset selection. §4 presents empirical results of our analyses and §5 discusses the implications for system design. We conclude in §6 with recommendations for further work. Our analysis work is released publicly to foster future research. [^1] https://github.com/cgshep/entropy-collapse-mobile-sensors
@@ 2.1 Mobile Sensor-based Security Applications
The notion of ‘hardness’ is typically inferred through model evaluation metrics. A general approach follows one whereby sensor data is collected from $N$ users from which various features are extracted in the time or frequency domain (e.g. cross-correlation, spectral energy, Hamming, Euclidean and mean-absolute distances) [4, 13, 3, 2]. Features are then classified using simple threshold-based or supervised classification models, e.g. Support Vector Machines (SVM) and Random Forests. Calculating false positive (FPR) and negative (FNR) rates [1], precision and recall [8], EERs [14, 13], Receiver-Operator Curves (ROC) [11], and accuracy [11] are used to evaluate the model with respect to distinguishing between legitimate and illegitimate samples. A system is deemed to be effective if the model can discriminate between such samples with low error. (User studies have also been employed to evaluate the effectiveness of sensor-based authentication systems [18, 1].)
Some proposals have attempted to assess the entropy of sensor signals within the context of a security mechanism, but this represents a minority of work in the literature. T2Pair by Li et al. [21], a zero-interaction pairing protocol, is found to have 32.3–38.5 bits of Shannon entropy. A refinement by Wu et al. [22] reported 51–54 bits of Shannon entropy. Even in the best cases, this is low relative to modern cryptographic standards. [^2] See AIS 20/31 [23]: FCS_RNG.1 specifies $\geq$240 and $\geq$250 bits of min- and Shannon entropy respectively for the effective internal state of a random number generator. We also point to adjacent work that has questioned the utility of mobile sensors. Markantonakis et al. [4], building on Gurulian et al. [5] and Shepherd et al. [6], presented a reproducibility study of mobile sensors when deployed under a 500ms time constraint for NFC-based transactions as specified by the EMV payment protocol. Here, 0.179–0.246 EER was reported depending on the given sensor combination. Sensors were thus deemed unsuitable for proximity and relay attack detection without posing usability and security issues in practice.
@@ 2.2 Sensor Entropy Analyses
In earlier work, Voris et al. [24] investigated accelerometers as true random number generators (TRNGs) on a WISP RFID tag and Nokia N97 phone. The authors find that min-entropy—defined in §2.3—is proportional to the motion applied to the device, with stationary movement having the lowest min-entropy. Intrinsic noise from the sensor’s circuitry and seismic noise, and the sampling rate of its analog-to-digital converter (ADC), are considered significant influences on entropy generation. Min-entropy values of 3.1–11.4 bits were measured depending on the movement of the accelerometer. Lv et al. [25] analysed three mobile sensors on an undisclosed Xiaomi Redmi smartphone: a triaxial accelerometer, gyroscope, and magnetometer. Min-entropy values of 0.593–5.876 are reported, depending on the modality and the entropy estimation method. Krhovják et al. [7] examined the entropy of image and audio data collected from mobile phone cameras and microphones respectively. Using Nokia N73 and E-Ten X500 and M700 phones, Shannon entropies of 2.9 (microphone) and 2.408–5.376 (camera) are reported, with min-entropy of 0.5 (microphone) and 0.754–3.928 (camera). Hennebert et al. [26] presented an analysis of 10 sensor modalities available on two wireless sensor monitors: a TI eZ430-RF2500 and a Zolertia Z1. A single-sensor analysis is presented, yielding min-entropy values of 0–7.85; motion sensors, e.g. accelerometer and vibration sensors, produced the highest entropy.
Sensors have also been suggested as entropy sources in low-cost RNG designs for mobile devices. Suciu et al. [27] proposed using a phone’s GPS module along with its accelerometer, gyroscope and orientation sensors. Using data from an HTC Google Nexus One, the approach passes the tests established in the NIST SP 800-22 [28] suite, but no precise entropy values were presented. Wallace et al. [29] explored an RNG design using the accelerometer, gyroscope, microphone, WiFi, GPS and camera data as randomness sources. Results are presented from a non-standard entropy evaluation using 37 Android devices.
@@
Sensors within existing work serve as inherently opportunistic entropy sources, i.e. in contrast with dedicated TRNGs using ring oscillators, Johnson-Nyquist thermal noise, and quantum phenomena (e.g. see [30]). Mobile sensors depend heavily on user behaviour and environmental, which could result in biases and correlations that are absent in controlled entropy sources. Existing sensor-based mechanisms largely overlook these dynamics, relying principally on heuristic or model evaluation metrics [8, 11]. Such approaches do not account well for skewed distributions and other biases in the underlying data that affect predictability. Contrast this with typical measures used in the area of randomness testing and authentication [31, 32, 30, 33, 23]. For instance, NIST SP800-90B [33] and AIS 20/31 [23] recommend the use of min-entropy to estimate ‘worst-case’ unpredictability of a given source. Our study bridges the gap by systematically evaluating entropy across modalities and datasets.
@@ 5.1 Comparison with Related Work
Prior studies have claimed or implied that mobile sensors are suitable data sources for security-critical applications. Much work has relied on model evaluation metrics as a proxy for evaluating such claims [13, 14, 4, 2], with a smaller subset using more established entropy metrics [22, 21]. However, analyses have shown that individual sensors confer very little entropy, especially under worst-case assumptions [25, 26, 24, 7]. Our results, derived from an analysis of multiple datasets and modalities (see Tables 25), corroborate and challenge these claims.